A 50 kg child climbs a 11.5-meter ladder to the top of a slide in 60 seconds.

a. As the boy slides down the slide and reaches the bottom, his speed is 12 m/s and the distance traveled along the slide is 108 m. Determine the average frictional force, in [N], acting on the boy?

b. Find the average power in [W].

c. As the child slides down the slide and reaches the bottom, his speed is 12 m/s and the distance traveled along the slide is 108 m. Find the mechanical energy, in [J], lost by the child in sliding down the slide.

d. Find the gained potential energy in [J].
1
Expert's answer
2020-08-04T16:00:22-0400

a. Let us calculate the potential energy of the boy when he is on the top of the slide ("g\approx 9.8\,\mathrm{N/kg}" ).

"E_p = mgh = 5635\,\mathrm{J}."

The final kinetic energy is

"E_k = \dfrac{mv^2}{2} = 3600\,\mathrm{J}."

The inequality of energies is due to the frictional force. The difference between the energies is

"\Delta E = E_p-E_k = 2035\,\mathrm{J}."

We calculate the average frictional force as

"F_f = \dfrac{\Delta E}{S} = \dfrac{2035\,\mathrm{J}}{108\,\mathrm{m}} \approx 18.8\,\mathrm{N}".

b. The average power while climbing is

"W = \dfrac{E_p}{\Delta t} = \dfrac{5635}{60} \approx 93.9\,\mathrm{W}."

c. The lost mechanical energy is "\Delta E = 2035\,\mathrm{J}."

d. The gained potential energy is "E_p = 5635\,\mathrm{J}."